What is the sum of the first 50 natural numbers?
Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

What is the sum of the first 50 natural numbers?
-----------
This is a arithmetic sequence.
Formula for the sum of n terms. S(n) = (n/2)(first + last)
Your Problem:
S(50) = (50)/2(1+50)
S(50) = 25*51 = 1275
==============
Cheers,
Stan H.

Answer by checkley75(3666)

the sum of the first 50 natural numbers is:
(n+1)*n/2=51*25=1275. or
n(n+1)/2=50(51)/2=2550/2=1275.